# Stability of Floquet sidebands and quantum coherence in 1D strongly interacting spinless fermions

### v2

@@ Abstract
A promising direction of research for realizing interesting quantum states of matter is Floquet engineering, where the interaction of the light field with electrons in materials [1, 2, 3, 4, 5, 6, 7] is used to tailor the band structure. For example, such periodically driven systems are predicted to offer a tunable platform to realize Fractional Chern insulators (FCIs) [8, 9], engineered topological states [10, 11, 12, 13, 14, 15], and coherent excitations in experiments with ultracold gases [16, 9, 17, 18, 19, 20] (for a review see Ref. 21). A hallmark of such periodically driven crystalline lattice structures is the emergence of Floquet-Bloch sidebands (FBs). Such FBs can directly be detected in pump–probe experiments by using time- and angle-resolved photoemission spectroscopy (trARPES) [22, 23, 24, 25, 26, 27, 28, 29, 30, 31], which gives insight into the time-dependent single-particle spectral function. The direct observation of FBs has been reported, e.g., for the topological insulator Bi2Se3 [23, 24], for the semiconductor material WSe2 [27], and recently for mono-layer graphene [28, 29] and for the topological antiferromagnet MnBi2Te4 [30]. These experimental findings are for weakly interacting systems. Model calculations and experimental work shows that it is non-trivial to understand the Floquet physics even for weak interactions [27, 32, 33]. This raises the question, whether for strongly interacting systems the same picture holds, or if the interplay of Floquet driving and many-body effects in the spectral function can modify the properties and stability of FBs [27].
Figure 1: Summary of the two extreme cases of our findings. (a) Sketch of model (1) and its phase diagram at half filling, with a quantum phase transition from a Luttinger liquid (LL) to a correlated charge density wave (CDW) insulator at $V/t_{h}=2$. (b) Typical result for the retarded spectral function $A^{\rm ret}_{k}(\omega)$ deep in the CDW insulating phase at $V/t_{h}=5$ at equilibrium. (c) and (d) Time evolution of the single particle propagator Eq. (4) in real space in the CDW phase for driving frequencies $\Omega=3.0$ and $\Omega=20.0$, respectively, at waiting time $t=0$, indicating the stability of quantum coherence in both regimes. (e) and (f) Time dependent spectral function $A^{\rm ret}_{k}(t,\omega)$ [Eq. (3)] in the CDW state at low and high frequency driving, $\Omega=3.0$ (no FBs visible) and $\Omega=20$ (FBs visible), respectively, at waiting time $t=10$. All results are at $V/t_{h}=5$.
Periodic driving leads to the dressing of electrons [34], energy absorption from photons [35, 36, 37], light-induced gaps [38, 39], and avoided crossings near the Fermi energy in the quasi-energy spectrum [15]. These effects have been explored using methods like Floquet dynamical mean-field theory (Floquet-DMFT) [40], which investigates non-equilibrium steady states. One finds band narrowing under strong fields, charge localization, and the coexistence of Wannier-Stark features with Floquet sidebands. Out-of-equilibrium DMFT [36, 37] and cluster perturbation theory [41] have also revealed photo-induced insulator-metal transitions. In strongly correlated systems with symmetry broken ground states, periodic driving can also induce in-gap features, such as Villain-type modes [42]. A major challenge in the presence of interactions in Floquet systems is the significant heating [43, 44, 45, 46, 47, 48, 49, 50, 51, 52], due to which a simple infinite-temperature state competes with Floquet engineering; due to this heating, one expects Floquet effects to play a role on some transient time scale, before the heating wins [44].
@@
All these aspects highlight how the choice of driving parameters can significantly alter the Floquet electronic structure of the system. The scope of this Letter is to investigate how for strongly correlated electron systems the interactions, the driving parameters, and noise influence the Floquet physics. We do so by directly studying the time evolution of the retarded time-dependent spectral function $A_{k}^{\rm ret}(t,\omega)$, of real space correlations, and of the momentum distribution function $\langle n_{k}\rangle(t)$, which relates to a Quantum Fisher Information (QFI) [53]. All results are obtained with time-dependent matrix product states (MPS) [54, 55] and exact diagonalization (ED) [56, 57, 58] techniques. We work on one-dimensional systems (1D), since MPS work best for these. This allows us to perform an unbiased study with high accuracy for the formation and stability of FBs and further Floquet effects in a typical strongly correlated system of 1D spinless fermions with nearest-neighbor interactions ($tV$-chain). Since in 1D correlation effects are strongest, this will help us in understanding the effect of interactions on the stability of FBs in correlated systems in general.
Recent inelastic neutron scattering studies on quantum magnetic materials show that it is insightful to complement the analysis of the spectral function by also studying the real-space time evolution of correlation functions and the QFI [59, 60, 61, 62]. Developments for the single-particle spectral function [62, 63] and out-of-equilibrium [64, 65] are under development. Here, we address the question, how coherence aspects of correlations in real space and the momentum distribution function, which is a simple QFI, behave in strongly interacting Floquet-driven systems.
@@
A summary showing the two extreme cases with and without FBs is sketched in Fig. 1. We find in the high frequency regime that the FBs are stable for long times even in the presence of strong interactions, and are only suppressed when adding noise to the drive. In the low frequency regime, they are suppressed if there is a large overlap between the FBs and the original spectral function. In general, our results for clean monochromatic driving indicate that the stability of the FBs depends on this overlap.
Model.—
@@
Here, $c_{j}^{(\dagger)}$ is the annihilation (creation) operator for a spinless fermion at site $j$, $n_{j}=c_{j}^{\dagger}c^{\phantom{\dagger}}_{j}$ is the operator of the local density on site $j$, $t_{h}$ the hopping strength between nearest neighboring lattice sites, and $V$ the strength of the interaction between particles sitting on neighboring sites. The periodic driving is realized as Peierls’ substitution [66] via a time-dependent vector potential $A_{V}(t)=A_{0}\sin(\Omega t)$, which is activated at time $t=0$. Unless stated otherwise, we use open boundary conditions (OBC) and the hopping parameter is set to $t_{h}\equiv 1$ throughout, as well as $\hbar\equiv 1$.
@@
Without the driving (i.e., for $A_{0}=0$), the model exhibits a Berezinskii-Kosterlitz-Thouless-like (BKT) phase transition from a Luttinger liquid (LL) [67] to a strongly correlated charge density wave (CDW) insulator at $V/t_{h}=2$, as obtained from Bethe ansatz after mapping to an XXZ spin-chain via Jorgan-Wigner transform [68]. In order to suppress artifacts due to the OBC, we apply a pinning-field $H_{\rm pin}=\mu n_{j}$ at one of the edge sites, which enforces the numerics to converge to one of the two possible CDW ground states [42, 69]. We compute the non-equilibrium generalization of the spectral function by performing a Fourier transform of the single-particle Green’s function,
$G_{\alpha\beta}^{\mathrm{ret}}\left(t,t^{\prime}\right)$ $=\theta\left(t-t^{\prime}\right)\left(G_{\alpha\beta}^{>}\left(t,t^{\prime}\right)-G_{\alpha\beta}^{<}\left(t,t^{\prime}\right)\right)$ (2)
@@
$=-i\theta\left(t-t^{\prime}\right)\left(\left\langle c_{\alpha}(t)c_{\beta}^{\dagger}\left(t^{\prime}\right)\right\rangle+\left\langle c_{\beta}^{\dagger}\left(t^{\prime}\right)c_{\alpha}(t)\right\rangle\right).$
@@
In non-equilibrium scenarios, one has to treat both time variables in Eq. (2) explicitly, which leads to a non-uniqueness of the Fourier transform (FT) to frequency space $\omega$ [70]. Typical choices are so-called Wigner coordinates, or relative time coordinates. The qualitative behavior typically does not depend on the particular choice of these coordinates [70, 71, 72]. In the following we choose a relative time coordinate $\tau=t-t^{\prime}$ for the FT, which simplifies the numerical approach by reducing the number of MPS states that need to be stored in the course of the calculation. In this way, we can interpret the evolution time $t$ as a waiting time, after which the spectral function is obtained under the influence of the periodic drive [69, 42]. The retarded single-particle spectral function then is computed via
$A_{k}^{\mathrm{ret}}(t,\omega)=-\operatorname{Im}\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\mathrm{d}\tau e^{i\omega\tau}W(\tau)G_{kk}^{\mathrm{ret}}(t,\tau)\,,$ (3)
where $W(\tau)$ is a windowing function. Typically, we use an adapted Tukey-window [73], since we find that for the weak signals of the Floquet side bands this gives the best resolution. The retarded spectral function Eq. (3) can then be interpreted as the spectral function measured at waiting time $t$, with the driving being turned on at $t=0$. In contrast to the equilibrium case, $A^{\rm ret}_{k}(t,\omega)$ can have negative values, see, e.g., Refs. 74, 75, which, however, in some strongly interacting systems have been found to be absent at later times [42, 75]. Here, in the context of the existence of FBs, we focus our discussion on the absolute value $|A^{\rm ret}_{k}(t,\omega)|$. Note that even at waiting time $t=0$ the results differ from the ground state behavior, since the Fourier-transform to $\omega$-space is done with the time-dependent Hamiltonian (1). For details of the MPS calculations and an error estimate, see the supplemental material.
Results.—
We define the spectral width $W$ for the ground state spectral function as the width of the interval $[\omega_{\rm min},\omega_{\rm max}]$ in which $A^{\rm ret}_{k}(\omega)\neq 0$, see Fig. 2(a). We identify $\Omega/t_{h}<W$ as a low-frequency drive and $\Omega/t_{h}>W$ as a high-frequency drive. In Fig. 2(a) we present MPS results for $A_{k}^{\rm ret}(\omega)$ at equilibrium deep in the CDW-regime for $L=32$ at half filling for $V/t_{h}=5$. Two continua of holon excitations, separated by a gap, are observed within the energy range $\omega\in[-10,10]$, consistent with previous studies [42]. The gap size $\Delta$ can be calculated using the Bethe ansatz, which for $V/t_{h}=5$ obtains $\Delta/t_{h}\approx 1.576$ [68]. This is in agreement with the gap size seen in Fig. 2(a), which is affected by finite-size effects and the limited spectral resolution of our methods. Fig. 2(d) shows ED results for $A_{k}^{\rm ret}(\omega)$ for $L=18$ at half filling for $V/t_{h}=1.5$. Figures 2(a) and (d) serve as the starting point for our investigation. A key aspect to consider is the potential overlap of the FBs with the original GS spectral function: If the excitation-continua of the replica and of the original spectral function overlap, interference or scattering events seem likely, so that the dynamics (in particular at later times) should be very different from the one in the high-frequency regime, where this overlap is not happening. This can be contrasted to non-interacting systems, where even in the presence of a complicated band structure (e.g., multiple bands), the absence of scattering terms should not lead to heating.
Figure 2: Retarded single-particle spectral function $A^{\rm ret}_{k}(t,\omega)$ [Eq. (3)] for the $tV$-chain (1). (a)-(c): $V/t_{h}=5$ for $L=32$ and OBC at half filling obtained with MPS. (a) Results at equilibrium. $W$ denotes the spectral width, in which $A^{\rm ret}_{k}(\omega)\neq 0$. (b) Driven case with $\Omega=3$ and $A_{0}=1.0$ at waiting time $t=0$. (c) The same driven case at waiting time $t=20$. The inset in (c) shows the result when keeping $k=\pi/2$ fixed. (d)-(f): $V/t_{h}=1.5$ for $L=18$ and PBC at half filling obtained with ED. (d) Results at equilibrium, $W^{\prime}$ denotes spectral width for $A^{\rm ret}_{k}(\omega)\neq 0$. (e) and (f) for the same driving parameters as in (b) and (c).
Low frequency drive.—
Floquet theory predicts that the FBs should be separated by integer($\pm$) multiples of the driving frequency $\Omega$. Additionally, periodic driving results in the dressing of electrons, which reduces the hopping strength to $t_{h}^{\rm eff}=t_{h}J_{0}(A_{0})$, with $J_{0}$ the 0th Bessel function, and where $A_{0}$ represents the amplitude of the drive [34, 35]. This expectation is clearly demonstrated in the non-interacting case [44]. To highlight the challenges associated with the formation of FBs in the presence of interactions, we fix the driving amplitude at $A_{0}=1.0$ and first choose a driving frequency $\Omega/t_{h}=3.0<W$ (low frequency drive). The results in CDW phase with $V/t_{h}=5.0$ are shown in Figs. 2(a)-(c). One would expect replicas of the entire equilibrium spectral function, which, due to the small value of $\Omega/t_{h}$, would significantly overlap with the original signal and with each other. In Fig. 2(b) we observe a significant spectral weight within the gap region, but no signature for FBs. The in-gap feature at $t=0$ exhibits a rich structure, which is reminiscent to the avoided-crossing scenario expected in non-interacting two-band models [36, 37, 15, 76, 32]. However, as shown in Fig. 2(c) for $t=20.0$, these features vanish at later times, and a broad continuum is obtained, with a stronger weight inside the original gap region. This melting of the gap features hints towards strong heating effects due to the chosen drive parameters, see further below. This indicates that once the replicas overlap in this strongly interacting system, the spectral function does not show FBs. Instead, strong heating is realized, which presumably is due to enhanced scattering of the particles caused by the overlapping FBs.
In comparison, in the gapless LL phase with $V/t_{h}=1.5$, also at $\Omega=3$ we obtain FBs, which, however, decay at long times, see Figs. 2(d)-(f), which were obtained using Lanczos time evolution, as the low-frequency case exhibits higher entanglement buildup in MPS, leading to lower accuracy (see the supp.mat for further discussion). Fig. 2(d) shows the spectral function in equilibrium, which follows a cosine-like dispersion with a buildup of a weakly populated excitation continuum without gap. As depicted in Fig. 2(d), we denote the spectral width by $W^{\prime}$ for which $A^{\rm ret}_{k}(\omega)\neq 0$. In Fig. 2(e), we observe an overlap of this weakly populated excitation continuum and the FBs at $t=0$. The FBs are clearly visible, even at $t=20$ in Fig. 2(f). However, the overlap of the FBs and the original spectral function now leads to a weaker signal of the FBs compared to the $t=0$ case (see the supp.mat for additional intermediate and longer waiting time instances). In addition, the intensity of the FBs is time dependent. Hence, our results indicate that in both the strongly interacting LL and CDW phases, overlapping bands are the main cause for the suppression of FBs in cleanly driven systems without noise, due to the enhancement of scattering and the resulting heating. However, if the overlap is weak as in the LL case, the FBs persist on a transient time scale, which for our parameters is $t_{\rm transient}\sim 50/t_{h}$ (see supp.mat.).
High frequency drive.—
In Fig. 3 we present results for the same set up, but with $A_{0}=1.6$, $\Omega/t_{h}=20>W$, and $V/t_{h}=5.0$.
Figure 3: The same as in Fig. 2 for $V/t_{h}=5$, but at high-frequency driving $\Omega=20$ and driving amplitude $A_{0}=1.6$. (a) and (b) show MPS results for $L=32$ and OBC, while the long-time results in (c) and (d) are obtained using Lanczos time evolution with $L=18$ sites and PBC. The inset in (a) shows the result when keeping $k=\pi/2$ fixed; the green highlighted region indicates the spectral function in the 0th Floquet sector, the blue shaded regions indicate the 1st FBs, and the magenta shaded regions indicate the 2nd FBs.
At waiting time $t=0.0$ in Fig. 3(a), we observe clear FBs at integer multiples of ±$\pm\Omega$. These FBs exhibit the full excitation continuum with renormalized bandwidth (since $t_{h}^{\rm eff}=J_{0}(A_{0})t_{h}$) and the emergence of an in-gap band, which was discussed in Ref. 42. This shows that also in the strongly interacting case the FBs are replicas of the full spectral function of the 0th Floquet sector, including additional features like the renormalization of the spectral width due to the dressed hopping $t_{h}^{\rm eff}$ and the in-gap signal caused by the driving. The question arises, if these FBs are stable in time, or if scattering between the particles eventually will destroy coherence, leading to a suppression of the FBs in $A^{\rm ret}_{k}(t,\omega)$ at later times. If scattering should be the reason for such a suppression, then a naive and rough estimate for the typical time scale would be $\sim 1/V$, which for our strongly interacting case would lead to a suppression of the FBs at waiting times $t\lesssim 1$. However, in Fig. 3(b), our results indicate that also at waiting time $t=20$, where for the low frequency drive all coherence was lost due to heating, the FBs remain stable, along with the in-gap band. We therefore numerically further test for the stability of the FBs in this strongly correlated case for much longer times, which are still amenable to numerical approaches. In order to increase the possible time scales, we employ the Lanczos time evolution method, but have to restrict to smaller system sizes of $L=18$ and use periodic boundary conditions (PBC). Figs. 3(c) and (d) show results for $t=100$ and $t=200$. Even at these substantially longer times, we can still resolve the FBs alongside the in-gap band. This demonstrates that high-frequency driving can yield FBs even in strongly interacting quantum many-body systems, exhibiting stability over the extended timescales we examined.
Stability of the FBs against noise—
The results so far indicate that in the high frequency regime stable FBs are obtained, as long as we have a coherent driving frequency. We now test this aspect by explicitly breaking the coherence by adding incoherent noise to the drive and ask for the stability of the FBs as a function of time. In order to reach longer times, we again apply the Lanczos time evolution method for $L=18$ (PBC) at half-filling, $V/t_{h}=5.0$, $A_{0}=1.6$, and a frequency $\Omega=20\pm\Omega_{noise}$, where $\Omega_{noise}$$\in[0,0.1]$ represents the random time-dependent noise in the driving frequency, such that $A_{V}(t)=A_{0}\sin((\Omega\pm\Omega_{noise})t)$. More precisely, we discretize the time variable such that $t=ndt$, with $n\in\mathbb{N}_{0}$, and then compute $A^{\rm ret}_{k}(t,\omega)$ when applying $A_{V}(t)=A_{0}\sin((\Omega\pm\Omega_{noise})ndt)$. Note that the results depend on the value of $dt$, as discussed in the supplemental material. Here, we present results for $dt=0.005$ in Fig. 4. In particular, Figs. 4(a), (b) and (d) demonstrate that, at small to intermediate waiting times, despite the introduction of noise, FBs are stable. However, at $t=20$ in Fig. 4(b), we observe a noticeable suppression of the second-order FBs. In Fig. 4(c), for waiting time $t=100$, this suppression becomes even more pronounced, with the cross-section at $k=\pi/2$ shown in Fig. 4(e) clearly showing the suppression of all FBs. Additionally, we notice a higher intensity in the in-gap band, as we found for the low frequency drive.
Figure 4: The same as in Fig. 3, but in the presence of time-dependent noise in the driving, where $\Omega=20\pm\Omega_{\rm noise}$ and $\Omega_{\rm noise}\in[0,0.1]$. The results are obtained using Lanczos time evolution with $L=18$ and PBC. Panels (a)-(c) show $A_{k}^{\rm ret}(t,\omega)$ at waiting times $t=0$, $t=20$, and $t=100$, respectively. Panels (d) and (e) show results at fixed $k=\pi/2$ at waiting times $t=0$ and $t=100$, respectively. All results show an average over 8 realizations of the random noise.
Real space correlations, heating, and Quantum Fisher Information.—
Our results so far have established that—in the absence of noise—the FBs are a stable feature in our strongly interacting system up to longer times. The question arises, how other observables behave when driving this strongly interacting system. Here, we study the time evolution of correlation functions in real space and of the momentum distribution function $\langle n_{k}\rangle(t)$, accompanied by the behavior of the energy as a function of time. This helps us to monitor the heating, which is intrinsic to driven interacting systems [48, 46, 49, 50, 51, 52]. In the following we focus on the case $V/t_{h}=5.0$.
We apply our low and high frequency driving protocols in Figs. 5(a)-(c) to the imaginary part of the real-space single-particle propagator
$G_{r,L/2}(t,\tau)=\bra{\psi(t)}\left[c^{\dagger}_{r}(\tau),c_{L/2}\right]\ket{\psi(t)}\,.$ (4)
Fig. 5(a) shows the typical light-cone-like behavior expected for gapped ground states [77, 78, 79, 80, 81, 82, 83, 84]. In addition, one observes as a function of time and position a ‘checkerboard-like’ pattern, which is similar to the one observed for spinons in neutron scattering experiments in spin correlation functions [59]. Here, we expect this should translate to the presence of holon excitations.
In stark contrast to the equilibrium results shown in Fig. 5(a), we see in Fig. 5(b) that for low frequency driving already at $t=0$ the correlation function is quickly suppressed as a function of time. This demonstrates that in this case coherence is quickly lost, which suggests the absence of FBs in $A_{k}^{\rm ret}(t,\omega)$. However, for the high-frequency driving case shown in Fig. 5(c), again a light-cone is visible. Clearly, it is more narrow, which is due to the reduced hopping $t_{h}^{\rm eff}=J_{0}(A_{0})t_{h}$, which leads to a lower propagation velocity of the perturbations in the system. In the equilibrium case one observes ‘wakes’ [59], which are created inside the light cone. In the driven case, they are more pronounced at later times, and every second one seems to be suppressed. Also, the checkerboard pattern is very similar to the one at equilibrium, but with slightly slower oscillation frequency in time. This is also obtained at later waiting times, see supplemental material. This illustrates that coherence is maintained, suggesting that here long-lived FBs are possible, as indeed found in the time-dependent spectral functions discussed above.
In addition to the real space correlations, coherence effects can also be studied by investigation the momentum distribution function $\langle n_{k}\rangle(t)=\langle c_{k}^{\dagger}c_{k}^{\phantom{\dagger}}\rangle(t)$, with $c_{k}^{(\dagger)}=1/\sqrt{L}\sum_{r=1}^{L}e^{(-)ik\cdot r}c_{r}^{(\dagger)}$. This quantity is accessible via time-of-flight experiments in cold-gases setups [16]. In Fig. 5(d) we show several snapshots in time for $\langle n_{k}\rangle(t)$ for $\Omega=3.0$ and $\Omega=20.0$ with and without noise obtained with ED for $L=18$ and $V/t_{h}=5$. At equilibrium, we obtain the typical result for a strongly correlated insulator, which is a smooth function without particular features. This behavior persists in the high frequency regime (without noise) up to the longest times studied by us, here $t=100$. However, when turning on the noise, or in the low frequency regime, $\langle n_{k}\rangle(t)$ at later times is essentially featureless and $\langle n_{k}\rangle\approx const$ for all values of $k$. In the low frequency case, this happens already at short times $t\sim 5$, while in the case with noise it happens only at later times $t\gtrsim 30$. This featureless result for $\langle n_{k}\rangle$ is expected at very high or infinite temperatures, further showing that in these cases coherence gets lost.
The findings for $G_{r,L/2}(t,\tau)$ and for $\langle n_{k}\rangle(t)$ can be related to heating effects by comparing with the behavior of the energy gain as a function of time $\Delta E(t)=\langle H(t)\rangle-E_{\rm GS}$ (with $E_{\rm GS}$ the ground state energy), see Fig. 5(e). (Further results for $G_{r,L/2}(t,\tau)$, $\langle n_{k}\rangle(t)$, $E(t)$ for longer times and at later waiting times, as well as a discussion of the numerical errors, can be found in the supplemental material.) This has been discussed extensively in the literature [48, 46, 49, 50, 51, 52]. When comparing the results for coherent driving with $\Omega=3$ and $\Omega=20$, we see a clear difference in the long time behavior. While for $\Omega=20$ (for all times treated by us) the energy oscillates around a value, which is very different from the one at infinite temperature, the case with $\Omega=3$ approaches this limit quickly. This seems in agreement with the expected suppression of heating at high driving frequencies [48]. When adding noise, we find intermediate behavior: while at times $t\lesssim 25$ the results for the coherent driving are reproduced, at later times the result at infinite temperatures is approached. This is in agreement with the findings of Ref. 85. It is interesting that in the presence of noise a clean transient time window exists, in which the behavior is essentially the one of the coherent driving, as can also be seen in Figs. 4(a) and (b).
Figure 5: Results at $V/t_{h}=5$: (a)-(c) show the imaginary part of $G_{r,L/2}(t,\tau)$ [Eq. (4)] (MPS, $L=32$, OBC) for the equilibrium case, and for the driving parameters of Figs. 2 and 3 at waiting time $t=0$. (d) shows $\langle n_{k}\rangle(t)$, for low-frequency driving at $\Omega=3$, and for high-frequency driving at $\Omega=20$ with and without noise for $L=18$ and PBC. (e) shows the time evolution of the energy gain compared to the initial state (ground state energy $E_{\rm GS}$), $\Delta E(t)=\langle H(t)\rangle-E_{\rm GS}$ for different driving frequencies for $L=14$ and PBC.
Conclusions and Outlook.—
We report the formation of FBs in the time-dependent spectral function of strongly interacting fermions. At low-frequency driving, they persist on a transient time scale, which depends on details of the system, if the overlap of the FBs and the original spectral function is small. For large overlap, we find fast loss of coherence and very strong heating, which leads to a suppression of the FBs. At high-frequency driving, even for strong interactions, the FBs are long lived. All features, including the many-body excitation continuum and a Villain-type in-gap mode, are replicated by the FBs. These, however, are suppressed in the presence of incoherent driving (noise).

### v3

@@ I Introduction
A promising direction of research for realizing interesting quantum states of matter is Floquet engineering, where the interaction of the light field with electrons in materials [1, 2, 3, 4, 5, 6, 7] is used to tailor the band structure. For example, such periodically driven systems are predicted to offer a tunable platform to realize Fractional Chern insulators (FCIs) [8, 9], engineered topological states [10, 11, 12, 13, 14, 15], and coherent excitations in experiments with ultracold gases [16, 9, 17, 18, 19, 20] (for a review see Ref. 21). A hallmark of such periodically driven crystalline lattice structures is the emergence of Floquet-Bloch sidebands (FBs). These FBs are replica of the original band structure and appear at energies which are integer multiples of the driving frequency. They can directly be detected in pump–probe experiments by using time- and angle-resolved photoemission spectroscopy (trARPES) [22, 23, 24, 25, 26, 27, 28, 29, 30], which gives insight into the time-dependent single-particle spectral function. The direct observation of FBs has been reported, e.g., for the topological insulator Bi2Se3 [23, 24], for the semiconductor material WSe2 [26], and recently for mono-layer graphene [28, 27]and for the topological antiferromagnet MnBi2Te4 [29]. These experimental findings are for weakly interacting systems. Model calculations and experimental work shows that it is non-trivial to understand the Floquet physics even for weak interactions [26, 31, 32]. Theoretical studies of interacting systems have been performed using cluster perturbation theory, or in infinite spatial dimensions using Floquet-DMFT and out-of-equilibrium DMFT  [33, 34, 35, 36]. However, in 1D and 2D strongly interacting systems the interplay of Floquet driving and many-body effects on the spectral functions and on the stability of FBs is still a topic of ongoing investigations [26].
Figure 1: Summary of the two extreme cases of our findings. (a) Sketch of a typical equilibrium result for the retarded spectral function $A^{\rm ret}_{k}(\omega)$ in the charge density wave (CDW) insulating phase at half filling for model (1). We denote by $W$ the spectral width (see Sec. III), in which the spectral function has populated continua. (b) and (c) illustrate the general expectation for FBs in such a strongly correlated system: In both sketches, the magenta region illustrates the zeroth Floquet sector, which is located at the position of the original equilibrium spectral function. Note that, due to the driving, in comparison to the equilibrium result the shape can be modified [37], and additional effects (e.g., in-gap bands [38]) can come into appearance. This spectral function in the zeroth Floquet sector then is replicated at energies, which are multiples of the driving frequency $\pm n\Omega$. Here, we only sketch the 1st replicas at $\pm\Omega$ (light blue). In (b) we sketch the situation, when the driving frequency $\Omega<W$. This leads to substantial overlap of the different Floquet sectors. In contrast to this, in (c) we sketch the situation if $\Omega>W$, so that there is no overlap between the Floquet sectors. (d) and (e) illustrate the effect of the driving on correlations in real space. They display the time evolution of the single particle propagator Eq. (5) in real space in the CDW phase for driving frequencies $\Omega=3.0$ and $\Omega=20.0$, respectively, at waiting time $t=5$, indicating the stability of quantum coherence in both regimes. (f) and (g) show examples for the resulting time dependent spectral function $A^{\rm ret}_{k}(t,\omega)$ [Eq. (4)] in the CDW state at low and high frequency driving, $\Omega=3.0$ (no FBs visible) and $\Omega=20$ (FBs visible at $\pm n\Omega$ indicated by the arrows), respectively, at waiting time $t=5$. All results are at $V/t_{h}=5$.
Floquet systems have also been explored in the context of many-body localization (MBL) and thermalization [39, 40, 41, 42, 43]. In particular, by fine-tuning drive parameters or suppressing hopping terms, one can induce localization even in clean systems (dynamical localization [44, 37]), revealing rich regimes of Floquet MBL and prethermal dynamics. Other important effects of periodic driving are energy absorption [45, 34, 35], dressing of electrons [37], light induced gaps [46, 47], metal-insulator transitions [34, 35, 36], and in strongly correlated systems it can also induce in-gap features in the presence of symmetry broken ground states [38]. A major challenge in interacting Floquet systems is the significant heating [43, 48, 49, 50, 51, 52, 53, 54, 55], which drives the system toward an infinite-temperature state. As a result, one expects Floquet effects to play a role on some transient time scale, before the heating wins [48].
@@
The scope of this paper is to investigate how for strongly correlated electron systems the interactions, the driving parameters, and noise influence the Floquet physics. We do so by directly studying the time evolution of the retarded time-dependent spectral function $A_{k}^{\rm ret}(t,\omega)$, of real space correlations, and of the momentum distribution function $\langle n_{k}\rangle(t)$, which relates to a Quantum Fisher Information (QFI) [56]. All results are obtained with time-dependent matrix product states (MPS) [57, 58] and exact diagonalization (ED) [59, 60, 61] techniques. While most of the calculations can be done using ED, it is useful to compare the results for the spectral functions with MPS results, since we can treat larger systems and hence reach a better resolution in $k$-space there. Studying larger systems will also be beneficial when exploring the dynamics of real-space correlations (see Sec. III.4). We focus on one-dimensional (1D) systems, as MPS methods are particularly efficient in this setting. Since correlation effects are strongest in 1D, we believe that studying them in this context will provide a generic understanding that is likely to hold in higher dimensions as well. This, in turn, will help us understand how interactions affect the stability of FBs in correlated systems in general. We complement this by studying coherence in real space correlations and in the momentum distribution function, which is a simple QFI, in strongly interacting Floquet-driven systems. This is motivated by recent inelastic neutron scattering studies on quantum magnets, which analyze the time evolution of correlation functions in real space and of the QFI [62, 63, 64, 65]. Further work relating the single-particle spectral function [65, 66] in- and out-of-equilibrium [67, 68] to QFIs is ongoing.
@@
A summary showing the two extreme cases with and without FBs is sketched in Fig. 1. We find in the high frequency regime that the FBs are stable for long times even in the presence of strong interactions, and are only suppressed when adding noise to the drive. In the low frequency regime, they are suppressed if there is a large overlap between the FBs and the original spectral function. In general, our results for clean monochromatic driving indicate that for interacting systems the stability of the FBs depends on this overlap.
This paper is organized as follows. In Sec. II, we introduce the model used in our study. Our main findings are presented in Sec. III, where we discuss the system’s behavior under a low-frequency drive (Sec. III.1), a high-frequency drive (Sec. III.2), the stability of the FBs against noise (Sec. III.3), and the evolution of real-space correlations, the momentum distribution function, and heating (Sec. III.4). We summarize our conclusions and provide an outlook in Sec. IV. Several technical details and complementary results are provided in the appendices. Details on the numerical methods are provided in Apps. A and B, with App. B discussing details of the Fourier transform used, in particular the choice of a window function. Further supporting results for the main text and an error analysis is provided in Apps. CDE, and F.
II Model
@@
Here, $c_{j}^{(\dagger)}$ is the annihilation (creation) operator for a spinless fermion at site $j$, $n_{j}=c_{j}^{\dagger}c^{\phantom{\dagger}}_{j}$ is the operator of the local density on site $j$, $t_{h}$ the hopping strength between nearest neighboring lattice sites, and $V$ the strength of the interaction between particles sitting on neighboring sites. The periodic driving is realized as Peierls’ substitution [69] via a time-dependent vector potential $A_{V}(t)=A_{0}\sin(\Omega t)$, which is activated at time $t=0$. Unless stated otherwise, we use open boundary conditions (OBC) and the hopping parameter is set to $t_{h}\equiv 1$ throughout, as well as $\hbar\equiv 1$. Without the driving (i.e., for $A_{0}=0$), the model Eq. (1) at half filling exhibits a Berezinskii-Kosterlitz-Thouless-like (BKT) phase transition from a Luttinger liquid (LL) [70] to a strongly correlated charge density wave (CDW) insulator at $V/t_{h}=2$, as obtained from Bethe ansatz after mapping to an XXZ spin-chain via Jorgan-Wigner transform [71]. In order to suppress artifacts due to the OBC, we apply a pinning-field $H_{\rm pin}=\mu n_{j}$ at one of the edge sites, which enforces the numerics to converge to one of the two possible CDW ground states [38, 72]. Throughout the paper, the initial state is prepared as a ground state of the system at temperature $T=0$. We compute the non-equilibrium generalization of the spectral function by performing a Fourier transform (FT) of the single-particle Green’s function,
@@
$G_{\alpha\beta}^{\mathrm{ret}}\left(t,t^{\prime}\right)$ $=-i\theta\left(t-t^{\prime}\right)\left\{\left\langle c_{\alpha}(t),\,c_{\beta}^{\dagger}\left(t^{\prime}\right)\right\rangle\right\}\,.$ (2)
@@
The FT to $\omega$-space is not unique [73]. Typical choices are Wigner coordinates, or relative-time coordinates. The qualitative behavior typically does not depend on the particular choice of these coordinates [73, 74, 75]. For our numerical approach, relative-time coordinates $\tau=t^{\prime}-t$ for the FT are more convenient, since in this way one can reduce the number of MPS states that need to be stored in the course of the calculation. The protocol is: